Amos is conducting a survey to find how many hours the students at his school spend doing

chores at home. After interviewing 50 people, he finds the mean answer is 8.2 hours with a standard

deviation of 1.5. What is the 95% confidence interval for his results?


1
Expert's answer
2021-05-07T10:24:25-0400

Out of 100%, 50 people say is 8.2hours

100-95= 5%

X is unknown for is result

95=====> x

50======>8.2

Crossing the multypling

50x=95×8.2

x= 779/50

x = 15.58 ans, confidence interval for his result